) A company has been making snowboards for many years. The monthly overhead is $7,900 and each snowboard costs $386 in materials and labor to make. If these items are sold for $427

(a) How many snowboards must be sold to break even?
Answer:

(b) What is the profit in dollars if 75 snowboards are made and sold? (Enter value but omit dollar sign.)
Answer:
Answer #1

a)

monthly overhead = 7900

snowboard cost = 386

let x be the no of snowboard sold

so total cost = 7900+386*x

now, total revenue= 427 * x

now,

for breakeven

total cost = total revenue

7900 + 386x = 427x

x = 192.6 = 193

b)

total cost = 7900+ 386*75

= 36850

total revenue = 427*75 = 32025

net loss = 36850-32025 = 4825